OpenStudy (anonymous):

The product of (a − b)(a − b) is a perfect square trinomial. Sometimes Always Never

OpenStudy (anonymous):

Do you know how to multiply them?

OpenStudy (anonymous):

a^2-b^2

OpenStudy (anonymous):

\you gotta multiply each term by the other

OpenStudy (anonymous):

\[(2+6)(3+8)\] for ex for this you gotta do 2 times 3+2 times 8+ 6times 3+ 6 times 8

OpenStudy (anonymous):

Sorry, do you know whata perferct sqaure trinomial is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok well its a trinomail that can be written as \[(a+b)^2=a^2+2ab+b^2\] or \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

do you multiply two binomails together?

OpenStudy (anonymous):

know how to multiply*

OpenStudy (anonymous):

you add

OpenStudy (anonymous):

well ok if you have \[(a+b)(c+d)=ac+ad+bc+bd\] fill it ion for your (a+b)(a+b)

OpenStudy (anonymous):

so its ab, ba

OpenStudy (anonymous):

(a+b)(a+b)=aa+ab+ab+bb=a^2+(2)ab+b^2

OpenStudy (anonymous):

does that fit a perfect square trinomial?

OpenStudy (anonymous):

I don't know what a perfect square trinomial is that's why I asked. is it?